### v2

# The Global Sections of Chiral de Rham Complexes on Closed Complex Curves

Abstract.
The space of global sections of the chiral de Rham complex on any closed complex curve with genus $g\geq 2$ is calculated.
B. Song is supported by National Natural Science Foundation of China Grant #12171447.
1. Introduction
The chiral de Rham complex $\Omega_{X}^{ch}$ is a sheaf of vertex superalgebras introduced by Malikov, Schechtman and Vaintrob in [14], which can be defined on either smooth manifolds, complex analytic manifolds or nonsingular algebraic varieties. It is graded by conformal weight and its weight zero part coincides with the ordinary de Rham sheaf. This sheaf has attracted significant attention in both the physics and mathematics literature. For example, in [9], Kapustin has shown that if X is a Calabi-Yau manifold, the cohomology of this sheaf can be identified with the infinite-volume limit of the half-twisted sigma model defined by E. Witten; in [3], Borisov presented an approach to toric Mirror Symmetry by the chiral de Rham complex.
The space of global sections $\Gamma(X,\Omega_{X}^{ch})$ is always a vertex superalgebra, and it is known to have extra structure when $X$ is endowed with geometric structures. For example, when $X$ is Kähler or hyperkähler, it has an $\mathcal{N}=2$ structure or $\mathcal{N}=4$ structure, respectively [2]. Recently in [17, 11], Linshaw and the first auther gave a complete description of $\Gamma(X,\Omega_{X}^{ch})$ for any compact Ricci-flat Kähler manifold $X$. In [4], for any congruence subgroup $G\subset SL(2,\mathbb{R})$, Dai constructed a basis of the $G$-invariant global sections of the chiral de Rham complex on the upper half plane, which are holomorphic at the cusps. The vertex operations are determined by a modification of the Rankin-Cohen brackets of modular forms. Further in [5], it is shown that if the global sections are allowed to be meromorphic at cusp, the space of $G$-invariant global sections will be a simple vertex algebra. In this paper, we extend the method developed in [17] to closed local symmetric space and calculate the space of global sections of chiral de Rham complex on any closed curves with genus $g\geq 2$ (Theorem 4.6).
The paper is organized as follows: in secion 2, we introduce the chiral de Rham complex; in section 3, we introduce the method of calculating the global sections of chiral de Rham complex in [17] and apply the method to Hermitian locally symmetric spaces. In section 4, we calculate the space of global sections for any closed complex curve $X$ with genus $g\geq 2$.
2. Chiral de Rham Complex
2.1. Vertex algebras
In this paper, we will follow the formalism of vertex algebras developed in [8]. A vertex algebra is the data $({\mathcal{A}},Y,L_{-1},1)$. In this notation,
  1. (1)
    ${\mathcal{A}}$ is a $\mathbb{Z}_{2}$-graded vector space over $\mathbb{C}$. The $\mathbb{Z}_{2}$-grading is called parity, and $|a|$ denotes the parity of a homogeneous element $a\in{\mathcal{A}}$.
  2. (2)
    $Y$ is an even linear map
    $Y:{\mathcal{A}}\to\operatorname{End}({\mathcal{A}})[[z,z^{-1}]],\qquad Y(a)=a(z)=\sum_{n\in\mathbb{Z}}a_{(n)}z^{-n-1}.$
    Here $z$ is a formal variable and $a(z)$ is called the field corresponding to $a$.
  3. (3)
    $1\in{\mathcal{A}}$ is called the vacuum vector.
  4. (4)
    $L_{-1}$ is an even endomorphism of ${\mathcal{A}}$.
They satisfy the following axioms:
  • Vacuum axiom. $L_{-1}1=0$; $1(z)=Id$; for $a\in{\mathcal{A}}$, $n\geq 0$, $a_{(n)}1=0$ and $a_{(-1)}1=a$;
  • Translation invariance axiom. For $a\in{\mathcal{A}}$,
    $[L_{-1},Y(a)]=\partial a(z);$
  • Locality axiom. Let $z,w$ be formal variables. For homogeneous $a,b\in{\mathcal{A}}$,
    $(z-w)^{k}[a(z),b(w)]=0$
    for some $k\geq 0$, where $[a(z),b(w)]=a(z)b(w)-(-1)^{|a||b|}b(w)a(z)$.
For $a,b\in{\mathcal{A}}$, $n\in\mathbb{Z}_{\geq 0}$, $a_{(n)}b$ is their $n^{\text{th}}$ product and their operator product expansion (OPE) is
$a(z)b(w)\sim\sum_{n\geq 0}(a_{(n)}b)(w)(z-w)^{-n-1}.$
The Wick product of $a(z)$ and $b(z)$ is $:a(z)b(z):\ =(a_{(-1)}b)(z)$. The other negative products are given by
$:\partial^{n}a(z)b(z):\ =n!(a_{(-n-1)}b)(z).$
2.2. $\beta\gamma-bc$ system.
Let $V$ be a $d$-dimensional complex vector space. The $\beta\gamma$-system $\mathcal{S}(V)$ and $bc$-system ${\mathcal{E}}(V)$ were introduced in [6]. A $\beta\gamma-bc$ system $\mathcal{W}(V)=\mathcal{S}(V)\otimes\mathcal{E}(V)$ is a vertex algebra generated by even elements $\beta^{i}$, $\gamma^{i}$, and odd elements $b^{i}$, $c^{i}$, where $i$ ranges from $1$ to $d$. The only nontrivial OPEs among them are:
$\beta^{i}(z)\gamma^{j}(w)\sim\delta_{ij}(z-w)^{-1},\quad b^{i}(z)c^{j}(w)\sim\delta_{ij}(z-w)^{-1}.$
Let $\mathcal{W}_{+}(V)$ be the subalgebra of $\mathcal{W}(V)$ generated by $\beta^{i}$, $\alpha^{i}=\partial\gamma^{i}$, $b^{i}$ and $c^{i}$. It is a topological vertex algebra with the following four fields:
$Q(z)=\sum^{d}_{i=1}:\beta^{i}(z)c^{i}(z):,\quad L(z)=\sum^{d}_{i=1}(:\beta^{i}(z)\partial\gamma^{i}(z):-:b^{i}(z)\partial{c^{i}}(z):),$
$J(z)=-\sum^{d}_{i=1}:b^{i}(z)c^{i}(z):,\quad G(z)=\sum^{d}_{i=1}:b^{i}(z)\partial\gamma^{i}(z):.$
The subalgebra $\mathcal{W}_{+}(V)$ is graded with conformal weights $k$ and fermionic numbers $l$:
$\mathcal{W}_{+}(V)=\bigoplus_{k\in\mathbb{Z}_{\geq 0}}\bigoplus_{l\in\mathbb{Z}}\mathcal{W}_{+}(V)[k,l].$
Here $\mathcal{W}_{+}(V)[k,l]=\{A\in\mathcal{W}_{+}(V)|L_{(1)}A=kA,J_{(0)}A=lA\}$.
2.3. Chiral de Rham Algebra.
Let $X$ be a complex manifold of dimension $d$. For a local coordinate system $(U,\gamma^{1},\ldots,\gamma^{d})$ of $X$, $\Omega^{ch}_{X}(U)$ is given by:
$\Omega^{ch}_{X}(U):=\mathcal{W}(\mathbb{C}^{d})\otimes_{\mathbb{C}[\gamma^{1}_{(-1)},\ldots,\gamma^{d}_{(-1)}]}\mathcal{O}(U).$
Here $\gamma^{i}_{(-1)}$ acts on the analytic function ring $\mathcal{O}(U)$ by function multiplications.
Then $\Omega^{ch}_{X}(U)$ is a vertex algebra generated by $\beta^{i}(z)$, $b^{i}(z)$, $c^{i}(z)$, and $f(z)$, $f\in\mathcal{O}(U)$, with relations:
$:f(z)g(z):=fg(z),\quad f,g\in\mathcal{O}(U).$
The only nontrivial OPEs among these generators are:
$\beta^{i}(z)f(w)\sim\dfrac{\frac{\partial f}{\partial\gamma^{i}}(z)}{z-w},\quad b^{i}(z)c^{j}(w)\sim\dfrac{\delta_{ij}}{z-w}.$
Let $\tilde{\gamma}^{1},\ldots,\tilde{\gamma}^{d}$ be a new coordinate system on $U$ with the relations:
$\tilde{\gamma}^{i}=f^{i}(\gamma^{1},\dots,\gamma^{d}),\quad\gamma^{i}=g^{i}(\tilde{\gamma}^{1},\ldots,\tilde{\gamma}^{d}).$
The following coordinate change equations is considered in [14]:
$\partial\tilde{\gamma}^{i}(z)=\sum^{d}_{j=1}:\dfrac{\partial f^{i}}{\partial\gamma^{j}}\partial\gamma^{j}(z):,$
$\tilde{b}^{i}(z)=\sum^{d}_{j=1}:\dfrac{\partial g^{j}}{\partial\tilde{\gamma}^{i}}(f(\gamma))(z)b^{j}(z):,$
$\tilde{c}^{i}(z)=\sum^{d}_{j=1}:\dfrac{\partial f^{i}}{\partial\gamma^{j}}c^{j}(z):,$
$\tilde{\beta}^{i}(z)=\sum^{d}_{j=1}(:\dfrac{\partial g^{j}}{\partial\tilde{\gamma}^{i}}(f(\gamma))(z)\beta^{j}(z):+\sum^{d}_{k=1}::\dfrac{\partial}{\partial\gamma^{k}}(\dfrac{\partial g^{j}}{\partial\tilde{\gamma}^{i}}(f(\gamma)))(z)c^{k}(z):b^{j}(z):)$
and for a general complex manifold $X$, the sections $L(z)$, $G(z)$, $Q_{(0)}$ and $J_{(0)}$ are globally defined.
The complex $\Omega^{ch}_{X}$ is graded as a sheaf of vertex algebras with conformal weights $k$ and fermion numbers $l$:
$\Omega^{ch}_{X}=\bigoplus_{k\in\mathbb{Z}_{\geq 0}}\bigoplus_{l\in\mathbb{Z}}\Omega^{ch}_{X}[k,l].$
Here $\Omega^{ch}_{X}[k,l](U)=\{A\in\Omega^{ch}_{X}(U)|L_{(1)}A=kA,J_{(0)}A=lA\}$ for any open subset $U$ of $X$.
Using the above construction, we obtain a smooth version of the chiral de Rham complex $\Omega_{X}^{ch,sm}$ on $X$, by regarding $X$ as a smooth real manifold, and replacing $\mathcal{W}(\mathbb{C}^{d})$, the complex coordinate system and the space of analytic functions $\mathcal{O}(U)$ by $\mathcal{W}(\mathbb{C}^{2d})$, the smooth coordinate system and the space of smooth complex functions $C^{\infty}(U)$, respectively.
The sheaf $\Omega_{X}^{ch,sm}$ includes $\Omega_{X}^{ch}$ and its complex conjugate. The sheaf $\Omega_{X}^{ch,sm}$ is graded by:
$\Omega^{ch,sm}_{X}=\bigoplus_{k\geq 0}\bigoplus_{l}\Omega^{ch,sm}_{X}[k,l].$
Here $\Omega^{ch,sm}_{X}[k,l](U)=\{A\in\Omega^{ch,sm}_{X}(U)|\bar{L}_{(1)}A=kA,\bar{J}_{(0)}A=lA\}$ for any open subset $U$ of $X$. Let $\Omega^{ch,l}_{X}=\Omega^{ch,sm}_{X}[0,l]$ and $\Omega^{ch,*}_{X}=\bigoplus_{l=0}^{d}\Omega^{ch,sm}_{X}[0,l]$. The operator $\bar{\partial}$ is the restriction of $\bar{Q}_{(0)}$ on $\Omega^{ch,*}_{X}$.
The following results in [17] will be used in the calculation of the space of global sections of the chiral de Rham complex on closed curves later:
Theorem 2.1.
(i) The chiral Hodge cohomology $H^{*}(X,\Omega^{ch}_{X})$ is equal to the cohomology of the complex $(\Omega^{ch,*}_{X},\bar{\partial})$.
(ii) The chiral Hodge cohomology $H^{*}(X,\Omega^{ch}_{X})$ is a vertex algebra. In particular, $\Gamma(X,\Omega^{ch}_{X})=H^{0}(X,\Omega^{ch}_{X})$ is a sub vertex algebra of $H^{*}(X,\Omega^{ch}_{X})$.
3. Antiholomorphic vector bundles corresponding to chiral de Rham complexes
Let $X$ be a compact Kähler manifold with a Kähler form $h$. $\bar{T}^{*}X$ is the antiholomorphic cotangent bundle of $X$ and $\bar{T}^{*}_{x}X$ is its fiber on $x\in X$.
3.1. Vector bundles induced from chiral de Rham complex.
In this section we recall the construction of antiholomorphic vector bundles corresponding to chiral de Rham complexes in [17]. For a holomorphic local coordinate system ($U$, $\gamma^{1},\ldots,\gamma^{d}$), $h$ is locally expressed as $h|_{U}=\sum_{i,j}H_{ij}d\gamma^{i}\wedge d\bar{\gamma}^{j}$. Let $(H^{jk})$ be the inverse of $(H_{ij})$, i.e., $\sum_{j}H_{ij}H^{jk}=\delta_{ik}$. Define the generators:
$\bm{\Gamma}^{j}=\sum_{i}:H_{ij}\partial\gamma^{i}:,\quad\bm{c}^{j}=\sum_{i}:H_{ij}c^{i}:,\quad\bm{b}^{j}=\sum_{i}:H^{ji}b^{i}:,\quad\bm{B}^{j}=Q_{(0)}\bm{b}^{j}.$
The only nontrivial OPEs among them are:
$\bm{B}^{i}(z)\bm{\Gamma}^{j}(w)\sim\delta_{ij}(z-w)^{-2},~{}~{}\bm{b}^{i}(z)\bm{c}^{j}(w)\sim\delta_{ij}(z-w)^{-1}.$
The vertex algebra generated by these elements is denoted by $\mathcal{W}^{\gamma}_{+}$, and it is isomorphic to $\mathcal{W}_{+}(V)$ as a vertex algebra by mapping $\bm{B}^{i}$, $\bm{\Gamma}^{i}$, $\bm{b}^{i}$ and $\bm{c}^{i}$ to $\beta^{i}$, $\alpha^{i}$, $b^{i}$ and $c^{i}$, respectively. $\bm{b^{i}}$, $\bm{c}^{i}$ and $\bm{\Gamma}^{i}$ commute with smooth functions on $U$, and
$\bm{B}^{j}(z)f(w)\sim\dfrac{\sum_{i}H^{ji}\frac{\partial f}{\partial\gamma^{i}}}{z-w}.$
Let $\tilde{\gamma}^{1},\ldots,\tilde{\gamma}^{d}$ be the new holomorphic local coordinates on $\tilde{U}$ with the relations $\gamma^{i}=g_{i}(\tilde{\gamma})$ and $\tilde{\gamma}^{i}=f_{i}(\gamma)$, the Kähler form is expressed locally as $h|_{\tilde{U}}=\sum_{i,j}\tilde{H}_{ij}d\tilde{\gamma}^{i}\wedge d\tilde{\gamma}^{j}$ in the new coordinate system, where $(\tilde{H}^{jk})$ is the inverse of $(\tilde{H}_{ij})$. Let
$\tilde{\bm{\Gamma}}^{j}=\sum_{i}:\tilde{H}_{ij}\partial\tilde{\gamma}^{i}:,\quad\tilde{\bm{c}}^{j}=\sum_{i}:\tilde{H}_{ij}\tilde{c}^{i}:,\quad\tilde{\bm{b}}^{j}=\sum_{i}:\tilde{H}^{ji}\tilde{b}^{i}:,\quad\tilde{\bm{B}}^{j}=Q_{(0)}\tilde{\bm{b}}^{j}.$
Then the following relations hold on $U\cap\tilde{U}$:
(3.1) $\tilde{\bm{\Gamma}}^{j}=\sum_{i}:\overline{\frac{\partial g_{i}}{\partial\tilde{\gamma}^{j}}}\bm{\Gamma}^{i}:,\quad\tilde{\bm{c}}^{j}=\sum_{i}:\overline{\frac{\partial g_{i}}{\partial\tilde{\gamma}^{j}}}\bm{c}^{i}:,\quad\tilde{\bm{b}}^{j}=\sum_{i}:\overline{\frac{\partial f_{j}}{\partial\gamma^{i}}}\bm{b}^{i}:,\quad\tilde{\bm{B}}^{j}=\sum_{i}:\overline{\frac{\partial f_{j}}{\partial\gamma^{i}}}\bm{B}^{i}:.$
So under the change of coordinates, both $\bm{B}^{i}$ and $\bm{b}^{i}$ change like $d\overline{\gamma}^{i}$, and both $\bm{\Gamma}^{i}$ and $\bm{c}^{i}$ change like $\frac{\partial}{\partial\overline{\gamma}^{i}}$.
Let $SW^{\gamma}$ be the polynomial ring generated by $\bm{b}^{i}_{(n)}$, $\bm{c}^{i}_{(n)}$, $\bm{\Gamma}^{i}_{(n)}$, $\bm{B}^{i}_{(n)}$, $1\leq i\leq d$, $n<0$, and $S^{\gamma}$ be the set of monomials in $SW^{\gamma}$. Let $SW(\bar{T}^{*}X)$ be an antiholomorphic vector bundle with fibers $SW(\bar{T}_{x}^{*}X)\cong SW^{\gamma}$ on any $x\in U$.
The set of monomials $S^{\gamma}$ is graded by:
$S^{\gamma}[k,l]=\{A\in S^{\gamma}|[L_{(1)},A]=kA,[J_{(0)},A]=lA\},$
$S^{\gamma}[k,l,s]=\{A\in S^{\gamma}[k,l]|\text{the number of $\bm{B}$'s in $A$ $-$ the number of $\bm{\varGamma}$'s in $A$ $=$ $s$}\}.$
For any $A\in S^{\gamma}$,
[…]
since $J_{(0)}\mathbb{b}^{i}=-\mathbb{b}^{i}$, $J_{(0)}\mathbb{c}^{i}=\mathbb{c}^{i}$, and $J_{(0)}\mathbb{B}^{i}=J_{(0)}\bm{\Gamma}^{i}=0$.
For any $A\in S^{\gamma}[k,l,s]$, $k\geq$ the number of $\mathbf{B}^{\prime}$s in $A$ plus the number of $\mathbf{\Gamma}^{\prime}$s in $A$. So when $k<|s|$, $S^{\gamma}[k,l,s]=\emptyset$.
Let $SW^{\gamma}[k,l]$, $SW^{\gamma}[k,l,s]$ be the subspaces of $SW^{\gamma}$ spanned by $S^{\gamma}[k,l]$, $S^{\gamma}[k,l,s]$, respectively. Accordingly, the subbundles $SW(\bar{T}^{*}X)[k,l]$ and $SW(\bar{T}^{*}X)[k,l,s]$ of $SW(\bar{T}^{*}X)$ have fibers isomorphic to $SW^{\gamma}[k,l]$, $SW^{\gamma}[k,l,s]$, respectively on any $x\in U$.
3.2. Chern connection.
The Kähler form $h$ induces a Hermition metric on $SW(\bar{T}^{*}X)$. Let $\nabla=\nabla^{1,0}+\bar{\partial}^{\prime}$ be the Chern connection corresponding to the metric, and it satisfies $\nabla ab=(\nabla a)b+(-1)^{|a|}a\nabla b$ for any $a,b\in\Omega^{*,*}_{X}(X,SW(\bar{T}^{*}X))$.
Let $\theta_{ij}=-\sum_{k}H^{ik}\bar{\partial}^{\prime}H_{kj}$ be a (0, 1)-form, then:
(3.2) $\nabla^{1,0}\bm{B}^{i}_{(k)}=\nabla^{1,0}\bm{\Gamma}^{i}_{(k)}=\nabla^{1,0}\bm{b}^{i}_{(k)}=\nabla^{1,0}\bm{c}^{i}_{(k)}=0$
and
(3.3) ¯$\bar{\partial}^{\prime}\bm{B}^{i}_{(k)}$ $=\sum_{j}\theta_{ij}\bm{B}^{j}_{(k)},\bar{\partial}^{\prime}\bm{b}^{i}_{(k)}=\sum_{j}\theta_{ij}\bm{b}^{j}_{(k)},$
¯$\bar{\partial}^{\prime}\bm{\Gamma}^{i}_{(k)}$ $=-\sum_{j}\theta_{ji}\bm{\Gamma}^{j}_{(k)},\bar{\partial}^{\prime}\bm{c}^{i}_{(k)}=-\sum_{j}\theta_{ji}\bm{c}^{j}_{(k)}$
for $k<0$. Using these expressions, for any $a\in S^{\gamma}$, the $\nabla a$ can be expressed as $\nabla a=\sum_{a^{\prime}\in S^{\gamma}}\Theta_{a,a^{\prime}}a^{\prime}$, where $\Theta_{a,a^{\prime}}$’s are one forms, and they are nonzero only if $a^{\prime}$ has the same number of $\bm{B}$’s, $\bm{\Gamma}$’s, $\bm{b}$’s and $\bm{c}$’s as $a$.
3.3. A canonical isomorphism of sheaves.
Let $\bar{\mathcal{O}}(SW(\bar{T}^{*}X))$ be the sheaf of antiholomorphic sections of $SW(\bar{T}^{*}X)$.
Let $\Omega^{0,k}_{X}(SW(\bar{T}^{*}X))$ be the sheaf of smooth (0, k)-forms with values in $SW(\bar{T}^{*}X)$ and let $\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))=\bigoplus\Omega^{0,k}_{X}(SW(\bar{T}^{*}X))$.
For any open subset $U$ of $X$, construct a vector space homomorphism $I_{U}$ locally:
$I_{U}:\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(U)\cong SW^{\gamma}\otimes_{\mathbb{C}}\Omega^{0,*}_{X}(U)\rightarrow\Omega^{ch,*}_{X}(U),a\otimes f\mapsto af,$
then by equation (3.1), a homomorphism $I$ of sheaves:
$I:\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))\rightarrow\Omega_{X}^{ch,*},$
is constructed. It is proved in [17] that:
Theorem 3.1.
The homomorphism $I$ is an isomorphism of $\bar{\mathcal{O}}(SW(\bar{T}^{*}X))$ modules.
Let $\bar{D}=I^{*}(\bar{\partial})$, i.e.,
$\bar{D}a=I_{U}^{-1}(\bar{\partial}I_{U}(a)),$
for any $a\in\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(U)$. According to the above Theorem 3.1 and Theorem 2.1, we get the following theorem,
Theorem 3.2.
$H^{i}(X,\Omega^{ch}_{X})\cong H^{i}(\Omega^{ch,*}_{X},\bar{\partial})\cong H^{i}(\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(X),\bar{D}).$
In particular:
$\Gamma(X,\Omega^{ch}_{X})\cong H^{0}(\Omega^{ch,*}_{X},\bar{\partial})\cong H^{0}(\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(X),\bar{D}).$
So we can use $H^{0}(\Omega^{0,*}_{X}(SW(\bar{T}^{*}X))(X),\bar{D})$ to calculate the space of global sections of chiral de Rham complex on $X$.
By [17], the operator $\bar{D}$ can be written in the following form:
$\bar{D}=\bar{\partial}^{\prime}+\sum_{i=1}^{\infty}F_{i}.$
Here $F_{i}$, mapping $\Omega^{0,*}_{X}(SW(\bar{T}X)[k,l,m])$ to $\Omega^{0,*+1}_{X}(SW(\bar{T}X)[k,l,m-i])$, is a first differential operator. Locally, for $a\in SW^{\gamma}$, $f\in\Omega^{0,*}_{X}(U)$,
(3.4) $I_{U}(F_{n}(a\otimes f))=\sum_{i,j}\sum_{s\in\mathcal{Z},|s|=n}\frac{1}{s!}[::\varGamma^{s}\mathbf{c}^{i}:\mathbf{b}^{j}:_{(0)},a]f\theta(s+e_{i},j)$
$+\sum_{j}\sum_{s\in\mathcal{Z},|s|=n+1}\frac{1}{s!}[:\varGamma^{s}\mathbf{B}^{j}:_{(0)}-\varGamma^{s}_{(-1)}\mathbf{B}^{j}_{(0)},a]f\theta(s,j)+\sum_{j}\sum_{s\in\mathcal{Z},|s|=n}\frac{1}{s!}[\varGamma^{s}_{(-1)},a](\mathbf{B}_{(0)}^{j}f)\theta(s,j).$
Here $\mathcal{Z}=\mathbb{Z}_{\geq 0}^{d}$ and $e_{i}=(0,\cdots,0,1,0\cdots,0)\in\mathcal{Z}$. For $s=(s_{1},\cdots,s_{d})\in\mathcal{Z}$,
$|s|=\sum_{i=1}^{d}s_{i},\quad s!=\prod_{i=1}^{d}s_{i}!.$
For any symbol ‘$X$’,
$X^{s}=\prod_{i=1}^{d}(X^{i})^{s_{i}}.$
For $s\in\mathcal{Z}$, $\theta(s+e_{i},j)=\mathbf{B}_{(0)}^{s}\theta_{ij}$. Let
$\varGamma^{i}(z)=\sum_{j\neq 0}\frac{1}{-j}\mathbf{\Gamma}^{i}_{(j)}z^{-j}.$
3.4. Hermitian locally symmetric space
A Hermitian locally symmetric space is a connected complex manifold with a Hermitian structure such that its curvature (1,1)-form is flat. Assume that $X$ is a Hermitian locally symmetric space and $R(\cdot,\cdot):TX\to TX$ is its curvature (1,1)-form. Locally
$R(\cdot,\cdot)\frac{\partial}{\partial{\gamma^{j}}}=\bar{\partial}^{\prime}(\partial H_{jk}H^{ki})\frac{\partial}{\partial{\gamma^{i}}}=-\bar{\partial}^{\prime}\bar{\theta}_{ij}\frac{\partial}{\partial{\gamma^{i}}}=-\overline{\partial\theta_{ij}}\frac{\partial}{\partial{\gamma^{i}}}.$
So
(3.5) $R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}}=-\overline{\mathbf{B}^{k}_{(0)}\theta_{ij}}\frac{\partial}{\partial{\gamma^{i}}}.$
And
$\overline{\partial(\mathbf{B}^{k}_{(0)}\theta_{ij})}\frac{\partial}{\partial{\gamma^{i}}}=\bar{\partial}^{\prime}(\overline{\mathbf{B}^{k}_{(0)}\theta_{ij}}\frac{\partial}{\partial{\gamma^{i}}})=-\bar{\partial}^{\prime}(R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}})=-(\nabla_{\bar{\partial}^{\prime}}R)(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}}=0.$
Thus
(3.6) $\partial(\mathbf{B}^{k}_{(0)}\theta_{ij})=0.$
We have the following lemmas for Hermitian locally symmetric spaces,
Lemma 3.1.
If $X$ is a Hermitian locally symmetric space, then,
$F_{n}=0,\quad n\geq 3,$
So
$\bar{D}=\bar{\partial}^{\prime}+F_{1}+F_{2},$
Locally, for $a\in SW^{\gamma}$, $f\in\Omega^{0,*}_{X}(U)$,
(3.7) $I_{U}(F_{1}(a\otimes f))=-\sum_{i,j,k}[::\varGamma^{k}\mathbf{c}^{i}:\mathbf{b}^{j}:_{(0)},a]f\overline{\langle R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}},d\gamma^{i}\rangle}$
$-\sum_{i,j,k}\frac{1}{2}[::\varGamma^{k}\varGamma^{i}:\mathbf{B}^{j}:_{(0)}-:\varGamma^{k}\varGamma^{i}:_{(-1)}\mathbf{B}^{j}_{(0)},a]f\overline{\langle R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}},d\gamma^{i}\rangle};$
(3.8) $I_{U}(F_{2}(a\otimes f))=-\sum_{i,j,k}\frac{1}{2}[:\varGamma^{k}\varGamma^{i}:_{(-1)},a](\mathbf{B}_{(0)}^{j}f)\overline{\langle R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}},d\gamma^{i}\rangle}.$
[…]
Proof.
If $X$ is a Hermitian locally symmetric space, then $\theta(s,j)=0$ if $|s|\geq 3$ by equation (3.6), and $F_{n}=0$ by equation (3.4). From equation (3.4) and the fact that $\varGamma^{i}_{(-1)}=0$,
$I_{U}(F_{1}(a\otimes f))$ $=$ $\sum_{i,j,k}[::\Gamma^{k}\mathbf{c}^{i}:\mathbf{b}^{j}:_{(0)},a]f\mathbf{B}^{k}_{(0)}\theta_{ij}$
$+\sum_{i,j,k}\frac{1}{2}[::\varGamma^{k}\varGamma^{i}:\mathbf{B}^{j}:_{(0)}-:\varGamma^{k}\varGamma^{i}:_{(-1)}\mathbf{B}^{j}_{(0)},a]f\mathbf{B}^{k}_{(0)}\theta_{ij};$
$I_{U}(F_{2}(a\otimes f))$ $=$ $\sum_{i,j,k}\frac{1}{2}[:\varGamma^{k}\varGamma^{i}:_{(-1)},a](\mathbf{B}_{(0)}^{j}f)\mathbf{B}^{k}_{(0)}\theta_{ij}.$
By equation (3.5),
$\mathbf{B}^{k}_{(0)}\theta_{ij}=-\overline{\langle R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}},d\gamma^{i}\rangle}.$
Lemma 3.2.
If $X$ is a Hermitian locally symmetric space, then
$[\nabla,F_{1}]=0.$
Proof.
In the second summation of the right-hand side of the equation (3.7), the term containing $\mathbf{B}^{j}_{(0)}$ is subtracted. Now $\nabla$ commutes with
$\sum_{k}\varGamma^{k}_{(m)}\otimes H^{kl}\frac{\partial}{\partial{\gamma^{l}}},\quad\sum_{i}\mathbf{c}^{i}_{(m)}\otimes d\bar{\gamma}^{i},\quad\sum_{i}\mathbf{\varGamma}^{i}_{(m)}\otimes d\bar{\gamma}^{i},\quad\sum_{j}\mathbf{b}^{j}_{(m)}\otimes\frac{\partial}{\partial\bar{\gamma}^{j}}$
for any $m\in\mathbb{Z}$, and
$\sum_{j}\mathbf{B}^{j}_{(n)}\otimes\frac{\partial}{\partial\bar{\gamma}^{j}},\quad n\neq 0$
From the expression of $F_{1}$ in equation (3.7) and the fact that $\nabla R=0$, we can see that $[\nabla,F_{1}]=0$. ∎
Locally, let $\tilde{F}_{2}^{j}$ be the operator on $\Omega^{0,*}(SW(\bar{T}^{*}X))$ such that
$I_{U}(\tilde{F}_{2}^{j}(a\otimes f))=-\sum_{i,k}\frac{1}{2}[:\varGamma^{k}\varGamma^{i}:_{(-1)},a]f\overline{\langle R(\cdot,\overline{H^{kl}\frac{\partial}{\partial{\gamma^{l}}}})\frac{\partial}{\partial{\gamma^{j}}},d\gamma^{i}\rangle}.$
Then
$F_{2}=\sum_{j,l}\tilde{F}_{2}^{j}H^{jl}\nabla_{\frac{\partial}{\partial\gamma^{l}}}.$
Lemma 3.3.
If $X$ is a Hermitian locally symmetric space, then
$[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},F_{2}]=0,\quad[\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},F_{2}]=\tilde{F}_{2}^{j}\tilde{R}(H^{jl}\frac{\partial}{\partial\gamma^{l}},\frac{\partial}{\partial\bar{\gamma}^{i}}).$
Here $\tilde{R}=[\bar{\partial}^{\prime},\nabla^{1,0}]$ is the curvature of $SW(\bar{T}^{*}X)$.
Proof.
$\nabla$ commutes with
$\sum_{k}\varGamma^{k}_{(m)}\otimes H^{kl}\frac{\partial}{\partial{\gamma^{l}}},\quad\sum_{i}\mathbf{\varGamma}^{i}_{(m)}\otimes d\bar{\gamma}^{i},\quad\sum_{j}H^{jl}\frac{\partial}{\partial\gamma^{l}}\otimes\frac{\partial}{\partial\bar{\gamma}^{j}}.$
By definition of $\tilde{F}_{2}^{j}$ and $\nabla R=0$, $\nabla$ commutes with $\tilde{F}_{2}^{j}\otimes H^{jl}\frac{\partial}{\partial\gamma^{l}}.$ Since
$[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},\nabla^{1,0}]=0,$
we have
$[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},F_{2}]=[\sum_{m}H^{im}\nabla_{\frac{\partial}{\partial\gamma^{m}}},\sum_{j,l}\tilde{F}_{2}^{j}H^{jl}\nabla_{\frac{\partial}{\partial\gamma^{l}}}]=0.$
$[\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},F_{2}]=[\nabla_{\frac{\partial}{\partial\bar{\gamma}^{i}}},\sum_{j,l}\tilde{F}_{2}^{j}H^{jl}\nabla_{\frac{\partial}{\partial\gamma^{l}}}]=\tilde{F}_{2}^{j}\tilde{R}(H^{jl}\frac{\partial}{\partial\gamma^{l}},\frac{\partial}{\partial\bar{\gamma}^{i}}).$
When $X$ be a closed complex curve with genus $g\geq 2$. We use the upper half plane as a covering space $\tilde{X}$ of $X$, then a Kähler form $h$ on $X$ is locally given by: $h|_{U}=Hd\gamma\wedge d\bar{\gamma}=\frac{i}{2y^{2}}d\gamma\wedge d\bar{\gamma}$, where $y=\operatorname{Im}\gamma>0$ and $\gamma$ is a holomorphic local coordinate. The generators of $\mathcal{W}_{+}^{\gamma}$ are $\mathbf{\Gamma},\mathbf{c},\mathbf{b},\mathbf{B}$. In this case $X$ is a Hermitian locally symmetric space,
$\theta=-H^{-1}\bar{\partial}^{\prime}H=-\frac{2}{\gamma-\bar{\gamma}}d\bar{\gamma},\quad\mathbf{B}_{(0)}\theta=id\bar{\gamma}.$
Thus by Lemma 3.1, we have
Corollary 3.1.
If $X$ is a complex curve with genus $g\geq 2$,
$\bar{D}=\bar{\partial}^{\prime}+F_{1}+F_{2},$
The differential operators $F_{1}$ and $F_{2}$ are locally expressed as:
(3.9) $I_{U}(F_{1}(a\otimes f))=i[::\varGamma\bm{c}:\bm{b}:_{(0)},a]fd\bar{\gamma}+\frac{i}{2}[::\varGamma^{2}:\bm{B}:_{(0)}-\varGamma^{2}_{(-1)}\mathbf{B}_{(0)},a]fd\bar{\gamma},$
and
(3.10) $I_{U}(F_{2}(a\otimes f))$ $=[\varGamma^{2}_{(-1)},a](\operatorname{Im}\gamma)^{2}(\nabla_{\frac{\partial}{\partial\gamma}}f)d\bar{\gamma},$
respectively. Here $a\in SW^{\gamma}$ and $f\in\Omega^{0,*}_{X}(U)$.
Lemma 3.4.
If $X$ is a complex curve with genus $g\geq 2$,
$iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}F_{2}=F_{2}iH^{-1}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}.$
Proof.
By Lemma 3.3,
$H^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}F_{2}$ $=$ $H^{-1}\nabla_{\frac{\partial}{\partial\gamma}}(F_{2}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{F}_{2}\tilde{R}(H^{-1}\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))$
$=$ $F_{2}H^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{F}_{2}H^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\tilde{R}(H^{-1}\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})$
$=$ $F_{2}H^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+F_{2}\tilde{R}(H^{-1}\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})$
$=$ $F_{2}H^{-1}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}.$
Here $F_{2}=\tilde{F}_{2}H^{-1}\nabla_{\frac{\partial}{\partial{\gamma}}}$ and $\tilde{F}_{2}$ commutes with $\nabla_{\frac{\partial}{\partial\gamma}}$ as in the proof of Lemma 3.3. ∎
4. Global Sections on closed Complex Curves
In this section we use Theorem 3.2 to calculate the space of global sections of chiral de Rham complex on closed curves with genus $g\geq 2$.
4.1. The curvature form on $SW(\overline{T}^{*}X)$.
For the curvature form $R\in\Omega^{1,1}(\operatorname{End}(TX))$:
$R(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})=\nabla_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}-\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial{\bar{\gamma}}}}.$
Using the Hermitian metric on the upper half plane, a direct computation shows:
$-iH^{-1}R(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})\frac{\partial}{\partial\gamma}=-\frac{\partial}{\partial\gamma}.$
So by equation (3.1) we have:
(4.1) $-iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})A=-A,\quad\text{if }A=\bm{B}_{(k)}\text{ or }A=\bm{b}_{(k)}.$
And:
(4.2) $-iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})A=A,\quad\text{ if }A=\bm{\Gamma}_{(k)}\text{ or }A=\bm{c}_{(k)}.$
Then we obtain the following result:
Lemma 4.1.
The action of the curvature form $\tilde{R}$ on a section $a_{s}$ of $SW(\bar{T}^{*}(X))[k,l,s]$ is:
(4.3) $-iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})a_{s}=(l-s)a_{s}$
Proof.
We observe that:
[…]
and
$=\text{the number of $\bm{B}$'s}-\text{the number of $\bm{\Gamma}$'s}.$ ’s .
So the equation (4.3) follows from the equations (4.14.2) above and the fact that $\tilde{R}_{E\otimes F}=\tilde{R}_{E}\otimes I_{F}+I_{E}\otimes\tilde{R}_{F}$ and $\tilde{R}_{E^{*}}=-\tilde{R}_{E}^{*}$ for any complex vector bundles $E$, $F$, here $E^{*}$ is the dual bundle of $E$. ∎
4.2. The equation $\bar{D}a=0$.
A nonzero section $a\in H^{0}(\Omega^{0,*}_{X}(SW(\bar{T}^{*}X)[k,l])(X),\bar{D})$, is a nonzero smooth section $a$ of $SW(\bar{T}^{*}X)[k,l]$ with $\bar{D}a=0$. The section $a$ can be written as
$a=\sum_{i\geq 0}a_{s-i},$
where $a_{s-i}$ is a smooth section of $SW(\bar{T}^{*}X)[k,l,s-i]$ and $a_{s}$ is nonzero. The summation is finite since $SW(\bar{T}^{*}X)[k,l,s-i]=\mathbf{0}$ when $k<i-s$. By Corollary 3.1 , $\bar{D}a=0$ is divided into a series of equations bellow:
$SW(\bar{T}^{*}X)[k,l,s]:$$\bar{\partial}^{\prime}a_{s}=0;$
$SW(\bar{T}^{*}X)[k,l,s-1]:$$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0;$
$SW(\bar{T}^{*}X)[k,l,s-2]:$$\bar{\partial}^{\prime}a_{s-2}+F_{1}a_{s-1}+F_{2}a_{s}=0;$
$\ldots$
$SW(\bar{T}^{*}X)[k,l,s-i]:$$\bar{\partial}^{\prime}a_{s-i}+F_{1}a_{s-i+1}+F_{2}a_{s-i+2}=0;$
$\ldots$
Therefore, $a_{s}$ must be a nonzero holomorphic section of $SW(\bar{T}^{*}X)[k,l,s]$.
Case 1: $l-s<0$
The following vanishing theorem for holomorphic sections of Hermitian vector bundles is well known (For a reference, see Chapter III of [10]),
Theorem 4.1.
Let E be a Hermitian vector bundle over a compact Hermition manifold M. Let R be the curvature of E, and $\hat{K}$ be the mean curvature of E. If $\hat{K}$ is negative-definite everywhere on M, then E admits no nonzero holomorphic sections.
When $l-s<0$, by Lemma 4.1, the mean curvature of $SW(\bar{T}^{*}X)[k,l,s]$ is negative-definite everywhere on $X$, so by above Theorem 4.1, there is no nonzero holomorphic section $a_{s}$.
Case 2: $l-s>0$.
Let $a_{s}$ a holomorphic section of $SW(\bar{T}^{*}X)[k,l,s]$ with $l-s>0$. Let $\bar{\partial}^{\prime*}$ be the dual operator of $\bar{\partial}^{\prime}$, then ¯$\bar{\partial}^{\prime*}=-iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}$.
(4.4) $\bar{\partial}^{\prime}\bar{\partial}^{\prime*}$ $=-d\bar{\gamma}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}$
$=-id\bar{\gamma}H^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}$
$=-id\bar{\gamma}H^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})).$
Lemma 4.2.
Given a nonzero holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,l,s]$ with $l-s>0$, $a_{s-1}=\frac{-1}{l-s}\bar{\partial}^{\prime*}F_{1}a_{s}$ is a solution of the equation ¯$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$.
Proof.
We have
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s}$ $=-id\bar{\gamma}H^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s}$
$=-id\bar{\gamma}H^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}a_{s}+(l-s)F_{1}a_{s}$
$=(l-s)F_{1}a_{s}.$
The last equality follows from Lemma 3.2 and the fact that $a_{s}$ is a holomorphic section. So $a_{s-1}=\frac{-1}{l-s}\bar{\partial}^{\prime*}F_{1}a_{s}$ is a solution of the equation. ∎
Lemma 4.3.
Suppose $a_{s-1}$ is given as in Lemma 4.2, then $a_{s-2}=\frac{-1}{2l-2s+1}(\bar{\partial}^{\prime*}F_{1}a_{s-1}+\bar{\partial}^{\prime*}F_{2}a_{s})$ is a solution of the equation ¯$\bar{\partial}^{\prime}a_{s-2}+F_{1}a_{s-1}+F_{2}a_{s}=0$.
Proof.
Since ¯$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$, we have:
$\nabla_{\frac{\partial}{\partial\bar{\gamma}}}a_{s-1}=-\iota_{\partial\bar{\gamma}}F_{1}a_{s}.$
We compute:
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-1}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s-1}$
$=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}F_{1}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}a_{s-1}+(l-s+1)F_{1}a_{s-1}$
$=d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}F_{1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}a_{s}+(l-s+1)F_{1}a_{s-1}.$
Using the expression in Lemma 4.2, we obtain:
$F_{1}a_{s-1}$ $=d\bar{\gamma}\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}\frac{-1}{l-s}\bar{\partial}^{\prime*}F_{1}a_{s}$
$=\frac{1}{l-s}d\bar{\gamma}\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}F_{1}a_{s}$
$=\frac{1}{l-s}iH^{-1}d\bar{\gamma}\nabla_{\frac{\partial}{\partial\gamma}}\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}a_{s}.$
Therefore,
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-1}=(2l-2s+1)F_{1}a_{s-1}.$
For the term ¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$, using Lemma 3.4, we have:
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{2}a_{s}$
$=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}F_{2}a_{s}+(l-s+1)F_{2}a_{s}$
$=-F_{2}iH^{-1}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}a_{s}+(l-s+1)F_{2}a_{s}.$
Moreover,
$iH^{-1}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}a_{s}$ $=iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})a_{s}+iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}a_{s}$
$=iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})a_{s}$
$=-(l-s)a_{s}.$
So:
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s}$ $=(l-s)F_{2}a_{s}+(l-s+1)F_{2}a_{s}$
$=(2l-2s+1)F_{2}a_{s}.$
Thus $a_{s-2}=\frac{-1}{2l-2s+1}(\bar{\partial}^{\prime*}F_{1}a_{s-1}+\bar{\partial}^{\prime*}F_{2}a_{s})$ is a solution of the equation. ∎
In general:
Theorem 4.2.
Let $a_{s+1}=0$, then the formula
(4.5) $a_{s-t}=\frac{-1}{t(l-s)+\frac{1}{2}t(t-1)}(\bar{\partial}^{\prime*}F_{1}a_{s-t+1}+\bar{\partial}^{\prime*}F_{2}a_{s-t+2})$
for $t\geq 1$ gives a solution of the equation $\bar{D}(\sum_{t\geq 0}a_{s-t})=0$.
Proof.
Induction on $t$. The cases $t=1,2$ are verified in Lemma 4.2 and Lemma 4.3, respectively. Next assume that $t\geq 3$.
By the inductive hypothesis ¯$\bar{\partial}^{\prime}a_{s-t+1}+F_{1}a_{s-t+2}+F_{2}a_{s-t+3}=0$, we have:
$\nabla_{\frac{\partial}{\partial\bar{\gamma}}}a_{s-t+1}+\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}a_{s-t+2}+\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{2}a_{s-t+3}=0.$
We compute:
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-t+1}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{1}a_{s-t+1}$
$=-iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}a_{s-t+1}+(l-s+t-1)F_{1}a_{s-t+1}$
$=-iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}F_{1}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}a_{s-t+1}+(l-s+t-1)F_{1}a_{s-t+1}$
$=iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}F_{1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(F_{1}a_{s-t+2}+F_{2}a_{s-t+3})+(l-s+t-1)F_{1}a_{s-t+1}.$
By the inductive formula of $a_{s-t+1}$,
$F_{1}a_{s-t+1}$ $=F_{1}\frac{-1}{(t-1)(l-s)+\frac{1}{2}(t-1)(t-2)}(\bar{\partial}^{\prime*}F_{1}a_{s-t+2}+\bar{\partial}^{\prime*}F_{2}a_{s-t+3})$
$=F_{1}\frac{1}{(t-1)(l-s)+\frac{1}{2}(t-1)(t-2)}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}(F_{1}a_{s-t+2}+F_{2}a_{s-t+3})$
$=\frac{1}{(t-1)(l-s)+\frac{1}{2}(t-1)(t-2)}iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}F_{1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(F_{1}a_{s-t+2}+F_{2}a_{s-t+3}).$
Therefore,
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s-t+1}$ $=((t-1)(l-s)+\frac{1}{2}(t-1)(t-2))F_{1}a_{s-t+1}+(l-s+t-1)F_{1}a_{s-t+1}$
$=(t(l-s)+\frac{1}{2}t(t-1))F_{1}a_{s-t+1}.$
Next we compute:
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s-t+2}$ $=-d\bar{\gamma}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}(\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}+\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}}))F_{2}a_{s-t+2}$
$=-iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}F_{2}a_{s-t+2}+(l-s+t-1)F_{2}a_{s-t+2}$
$=-F_{2}iH^{-1}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}a_{s-t+2}+(l-s+t-1)F_{2}a_{s-t+2}.$
Moreover, by the inductive hypothesis ¯$\bar{\partial}^{\prime}a_{s-t+2}+F_{1}a_{s-t+3}+F_{2}a_{s-t+4}=0$, we have:
$\nabla_{\frac{\partial}{\partial\bar{\gamma}}}a_{s-t+2}+\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}a_{s-t+3}+\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{2}a_{s-t+4}=0.$
So:
$~{}~{}~{}-iH^{-1}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}a_{s-t+2}$
$=-iH^{-1}\tilde{R}(\frac{\partial}{\partial\gamma},\frac{\partial}{\partial\bar{\gamma}})a_{s-t+2}-iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\nabla_{\frac{\partial}{\partial\bar{\gamma}}}a_{s-t+2}$
$=(l-s+t-2)a_{s-t+2}+iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{1}a_{s-t+3}+iH^{-1}\nabla_{\frac{\partial}{\partial\gamma}}\iota_{\frac{\partial}{\partial\bar{\gamma}}}F_{2}a_{s-t+4}.$
By the inductive formula of $a_{s-t+2}$,
$F_{2}a_{s-t+2}$ $=F_{2}\frac{-1}{(t-2)(l-s)+\frac{1}{2}(t-2)(t-3)}(\bar{\partial}^{\prime*}F_{1}a_{s-t+3}+\bar{\partial}^{\prime*}F_{2}a_{s-t+4})$
$=\frac{1}{(t-2)(l-s)+\frac{1}{2}(t-2)(t-3)}F_{2}iH^{-1}\iota_{\frac{\partial}{\partial\bar{\gamma}}}\nabla_{\frac{\partial}{\partial\gamma}}(F_{1}a_{s-t+3}+F_{2}a_{s-t+4}).$
Therefore,
¯¯$~{}~{}~{}\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{2}a_{s-t+2}$
$=((l-s+t-2)+((t-2)(l-s)+\frac{1}{2}(t-2)(t-3))+(l-s+t-1))F_{2}a_{s-t+2}$
$=(t(l-s)+\frac{1}{2}t(t-1))F_{2}a_{s-t+2}.$
So under the inductive hypothesis, $a_{s-t}=\frac{-1}{t(l-s)+\frac{1}{2}t(t-1)}(\bar{\partial}^{\prime*}F_{1}a_{s-t+1}+\bar{\partial}^{\prime*}F_{2}a_{s-t+2})$ is a solution of the equation ¯$\bar{\partial}^{\prime}a_{s-t}+F_{1}a_{s-t+1}+F_{2}a_{s-t+2}=0$. ∎
So for any $a_{s}\in H^{0}(X,SW(\bar{T}^{*}X))[k,l,s]$ with $l-s>0$, we can give an $a=\sum_{i\geq 0}{a_{s-i}}$ satisfying the condition for global sections.
Case 3: $l-s=0$.
In this case, we assume that $a_{s}\in H^{0}(X,SW(\bar{T}^{*}X))[k,s,s]$.
Lemma 4.4.
$SW(\bar{T}^{*}X)[k,s,s]$ is a trivial vector bundle. So every holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,s,s]$ satisfies $\nabla a=0$ and the space of holomorphic sections of $SW(\bar{T}^{*}X)[k,s,s]$ is isomorphic to its fibre $SW^{\gamma}[k,s,s]$.
Proof.
By the construction of $SW(\bar{T}^{*}X)[k,l,s]$, it is a direct sum of tensors of $\bar{T}^{*}X$ and $\bar{T}X$ when $X$ is a curve. When $l=s$, the number of $\bar{T}^{*}X$ and $\bar{T}X$ are equal in the tensor. So $SW(\bar{T}^{*}X)[k,s,s]$ is a trivial vector bundle.
Corollary 4.1.
For any holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,s,s]$, we have $F_{2}\,a_{s}=0$.
Proof.
This follows directly from equation (3.10) and Lemma 4.4. ∎
The Hodge decomposition of forms will be used later in the calculation of global sections:
Theorem 4.3.
Let $X$ be a compact Hermitian manifold, and $E$ an Hermitian vector bundle over $X$, the operator $\Delta$ is given by $\Delta=\bar{\partial}^{\prime}\bar{\partial}^{\prime*}+\bar{\partial}^{\prime*}\bar{\partial}^{\prime}$, $\Omega_{X}^{p,q}(E)$ is the space of $(p,q)$-forms over $X$ taking values in $E$, then:
(4.6) $\Omega_{X}^{p,q}(E)=\operatorname{Ker}\Delta|_{\Omega_{X}^{p,q}(E)}\oplus\bar{\partial}^{\prime}\Omega_{X}^{p,q-1}(E)\oplus\bar{\partial}^{\prime*}\Omega_{X}^{p,q+1}(E).$
Lemma 4.5.
There exists a section $a_{s-1}$ of $SW(\bar{T}^{*}X)[k,s,s-1]$ such that ¯$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$ if and only if $F_{1}a_{s}=0$.
Proof.
As in the proof of Lemma 4.2, we have:
¯¯$\bar{\partial}^{\prime}\bar{\partial}^{\prime*}F_{1}a_{s}=(s-s)F_{1}a_{s}=0.$
On the other hand:
¯¯$\bar{\partial}^{\prime*}\bar{\partial}^{\prime}F_{1}a_{s}=0.$
Therefore,
$\Delta F_{1}a_{s}=0.$
So ¯$\bar{\partial}^{\prime}a_{s-1}+F_{1}a_{s}=0$, if and only if
$F_{1}a_{s}\in\operatorname{Ker}\Delta\cap\operatorname{im}\bar{\partial}^{\prime}.$
That is $F_{1}a_{s}=0$ by Theorem 4.3. ∎
Theorem 4.4.
For a holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,s,s]$, there is an $a=\sum_{i\geq 0}a_{s-i}$ with $\bar{D}a=0$ if and only if $F_{1}a_{s}=0$. If $F_{1}a_{s}=0$, then $\bar{D}a_{s}=0$.
Proof.
By Lemma 4.5, if there is an $a=\sum_{i\geq 0}a_{s-i}$ with $\bar{D}a=0$, we must have $F_{1}a_{s}=0$. By Corollary 3.1 and Corollary 4.1, if $F_{1}a_{s}=0$, then $\bar{D}a_{s}=0$. ∎
So a holomorphic section $a_{s}$ of $SW(\bar{T}^{*}X)[k,s,s]$ corresponds to a global section of $\Omega^{ch}_{X}$ if and only if $F_{1}a_{s}=0$ and in this case $a_{s}$ itself is a global section.
4.3. The $\mathfrak{sl}_{2}$ action on a $\beta\gamma-bc$ system
In this subsection, we discuss the equation $F_{1}a_{s}=0$ when $a_{s}$ is a holomorphic section of $SW(\bar{T}^{*}X)[k,s,s]$.
Let $V$ be a complex linear space of dimension one. Note that when the complex manifold $X$ is of dimension one, $\mathcal{W}^{\gamma}_{+}$ is isomorphic to $\mathcal{W}_{+}(V)$ as a vertex algebra by mapping $\bm{B}$, $\bm{\Gamma}$, $\bm{b}$, $\bm{c}$ to $\beta$, $\alpha$, $b$, $c$, respectively. Let $\tilde{\gamma}=\gamma-\gamma_{(-1)}=\sum_{n\neq-1}\gamma_{(n)}z^{-n-1}$, then the subalgebra of $\mathcal{W}_{+}(V)$ generated by $\bm{B}$, $\varGamma$, $\bm{b}$, $\bm{c}$ is isomorphic to the subalgebra of $\mathcal{W}_{+}$ generated by $\beta$, $\tilde{\gamma}$, $b$, $c$ by mapping $\bm{B}$, $\varGamma$, $\bm{b}$, $\bm{c}$ to $\beta$, $\tilde{\gamma}$, $b$, $c$, respectively.
Let $T$ be the Lie algebra generated by vector fields $\{x^{i}\frac{\partial}{\partial x}\}_{i=0,1,2}$ on $V$. $T$ is isomorphic to $\mathfrak{sl}_{2}$ with the following relations:
$[x^{i}\frac{\partial}{\partial x},x^{j}\frac{\partial}{\partial x}]=(j-i)x^{i+j-1}\frac{\partial}{\partial x}.$
Let $\mathcal{L}:T\rightarrow End_{\mathbb{C}}(\mathcal{W}(V))$ be the Lie algebra homomorphism (See Part III in [13]) given by
$\mathcal{L}(x^{i}\frac{\partial}{\partial x})=(Q_{(0)}:\gamma^{i}b:)_{(0)},$
where $Q=:\beta c:$.
Therefore, $\mathcal{L}$ is an action of $\mathfrak{sl}_{2}$ on a $\beta\gamma-bc$ system $\mathcal{W}(V)$. And by the isomorphism between $\mathcal{W}^{\gamma}_{+}$ and $\mathcal{W}_{+}(V)$, $\mathcal{L}$ also induces a natural action of $\mathfrak{sl}_{2}$ on $SW^{\gamma}$.
Define $\mathcal{L}^{+}:T\rightarrow End_{\mathbb{C}}(\mathcal{W}(V))$ by:
$\mathcal{L}^{+}(x^{i}\frac{\partial}{\partial x})=(:\tilde{\gamma}^{i}\beta:)_{(0)}-i(::\tilde{\gamma}^{i-1}b:c:)_{(0)}.$
Lemma 4.6.
For the map $\mathcal{L}$ and the map $\mathcal{L}^{+}$, the following relations hold:
$\mathcal{L}(\frac{\partial}{\partial x})$ $=$ $\mathcal{L}^{+}(\frac{\partial}{\partial x}),$
$\mathcal{L}(x\frac{\partial}{\partial x})$ $=$ $\mathcal{L}^{+}(x\frac{\partial}{\partial x})+\gamma_{(-1)}\mathcal{L}^{+}(\frac{\partial}{\partial x}),$
$\mathcal{L}(x^{2}\frac{\partial}{\partial x})$ $=$ $\mathcal{L}^{+}(x^{2}\frac{\partial}{\partial x})+\gamma_{(-1)}\mathcal{L}^{+}(2x\frac{\partial}{\partial x})+\gamma_{(-1)}^{2}\mathcal{L}^{+}(\frac{\partial}{\partial x}).$
Proof.
See Lemma 6 in [17]. ∎
Let $\mathcal{W}^{T}(V)$ be the space of invariants in $\mathcal{W}(V)$ under the action of $\mathfrak{sl}_{2}$:
Definition 4.1.
The subspace $\mathcal{W}^{T}(V)$ of $\mathcal{W}(V)$ is defined by
$\mathcal{W}^{T}(V)=\left\{A\in\mathcal{W}(V)|\mathcal{L}(x^{i}\frac{\partial}{\partial x})A=0,i=0,1,2\right\}.$
Theorem 4.5.
The subspace of holomorphic sections of $\bigoplus_{k,s}SW(\bar{T}^{*}X)[k,s,s]$ with $F_{1}a_{s}=0$ is isomorphic to $\mathcal{W}^{T}(V)$.
Proof.
By Lemma 4.4, the space of holomorphic sections of $\bigoplus_{k,s}SW(\bar{T}^{*}X)[k,s,s]$ is isomorphic to its fibre $\bigoplus_{k,s}SW^{\gamma}[k,s,s]$. That is the subspace of $SW^{\gamma}$ generated by monomials with the number of $\mathbf{B}$’s plus the number of $\mathbf{b}$’s is equal to the number of $\mathbf{\Gamma}$’s plus the number of $\mathbf{c}$’s. By Lemma 4.6,
$\mathcal{W}^{T}(V)=\{A\in\mathcal{W}(V)|\mathcal{L}^{+}(x^{i}\frac{\partial}{\partial x})A=0,i=0,1,2\}.$
Now
$\mathcal{L}^{+}(\frac{\partial}{\partial x})A=\beta_{(0)}A=0,$
if and only if $A\in\mathcal{W}_{+}(V)$.
$\mathcal{L}^{+}(x\frac{\partial}{\partial x})A=:\tilde{\gamma}\beta:_{(0)}A-:bc:_{(0)}A=0,$
if and only if the number of $\beta$’s plus the number of $b$’s is equal to the number of $\alpha$’s plus the number of $c$’s in each monomial of $A$.
$\mathcal{L}^{+}(x^{2}\frac{\partial}{\partial x})A=:\tilde{\gamma}\tilde{\gamma}\beta:_{(0)}A-2:\tilde{\gamma}bc:_{(0)}A=0,$
is equivalent to $F_{1}a_{s}=0$ by the isomorphism between $SW^{\gamma}$ and $\mathcal{W}_{+}(V)$. ∎
4.4. The space of global sections.
We derive a description of the space of global sections of $\Omega^{ch}_{X}$ on the complex curve $X$ from the computations in the preceding part:
Theorem 4.6.
If $X$ is a closed complex curve with genus $g\geq 2$, then the space of global sections:
(4.7) $H^{0}(X,\Omega^{ch}_{X})\cong\bigoplus_{k,s}((H^{0}(X,SW(\bar{T}^{*}X))[k,s,s]\cap\operatorname{Ker}F_{1})\bigoplus(\bigoplus_{l>s}H^{0}(X,SW(\bar{T}^{*}X))[k,l,s]))$
Moreover, the subspace $\bigoplus_{k,s}(H^{0}(X,SW(\bar{T}^{*}X))[k,s,s]\cap\operatorname{Ker}F_{1})$ is isomorphic to the space of invariants under the action of $\mathfrak{sl}_{2}$ on the $\beta\gamma-bc$ system $\mathcal{W}(V)$, where $V$ is a complex linear space of dimension one.